OpenStudy (warpedkitten):

Zack, Rachel, and Maddie are unraveling a huge ball of yarn to see how long it is. As they move away from each other, they form a triangle. The distance from Zack to Rachel is 3 meters. The distance from Rachel to Maddie is 2.5 meters. The distance from Maddie to Zack is 4 meters. Find the measures of the three angles in the triangle.

OpenStudy (anonymous):

looks like a law of cosines problem to me

OpenStudy (warpedkitten):

It is, but I'm not understanding it. I've done similar problems and it seems like every time I try, I don't get any of the answer choices.

OpenStudy (anonymous):

lets agree on this labelling, i will let you fill in the names etc

OpenStudy (anonymous):



OpenStudy (anonymous):

B is opposite the largest side, so it is the largest angle, we can find that first

OpenStudy (anonymous):

\[b^2=a^2+c^2-2ac\cos(B)\] but we want the angle so \[\cos(B)=\frac{a^2+c^2-b^2}{2ac}\]

OpenStudy (anonymous):

filling in the numbers we get \[\cos(B)=\frac{3^2+2.5^2-4^2}{2\times 3\times 2.5}\]

OpenStudy (anonymous):

take the arccosine to find B

OpenStudy (princeevee):

satellite can you help me?

OpenStudy (anonymous):

i must have made some mistake here

OpenStudy (anonymous):

oh, no i didn't here use the calculator http://www.wolframalpha.com/input/?i=arccos%28%283^2%2B2.5^2-4^2%29%2F%282*3*2.5%29%29

OpenStudy (anonymous):

looks like the largest angle is \(92.87^\circ\) find the next one of your choice using the law of sines

OpenStudy (warpedkitten):

I've got it now, thank you!